Ask your own question, for FREE!
Mathematics 7 Online
OpenStudy (anonymous):

A container contains red, white, and yellow balls. Suppose that once a ball is drawn it is not replaced. What is the probability that red balls will be drawn in tries? (Express your answer as a fraction.)

OpenStudy (anonymous):

I think your Numlock key acted up

OpenStudy (anonymous):

A container contains 2 red, 3 white, and 5 yellow balls. Suppose that once a ball is drawn it is not replaced. What is the probability that 2 red balls will be drawn in 2 tries? (Express your answer as a fraction.)

OpenStudy (anonymous):

2 red out of 10 possibilities, then 1 out of 9. Mathematically, that's 2/10*1/9 = 1/45

OpenStudy (savvy):

(2/10)X(1/9) = 1/45

Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!
Latest Questions
Arriyanalol: i need help
2 days ago 7 Replies 1 Medal
Arriyanalol: CONGRATS to me
2 days ago 8 Replies 2 Medals
Legomaster346: What colleges are good that are near Arab Alabama?
1 week ago 5 Replies 0 Medals
Arriyanalol: i need help
1 week ago 6 Replies 2 Medals
Can't find your answer? Make a FREE account and ask your own questions, OR help others and earn volunteer hours!

Join our real-time social learning platform and learn together with your friends!